# Asymptotic Behavior of Integral Projection Models via Genealogical Quantities

### v2

@@ 1. Introduction
The Integral Projection Model (IPM) is a mathematical framework for describing transitions in populations with continuous state variables (Ellner and Rees, 2006). In discrete-time settings, IPMs have been developed and applied widely in ecology and demography. For example, Coulson et al. (2010) studied trait evolution, while Coulson and Tuljapurkar (2008) analyzed asymptotic population dynamics under environmental variability. These and other studies (White et al., 2016; Nicole et al., 2011; Doak et al., 2021; Merow et al., 2014; Elderd and Miller, 2016) have linked spectral theory to equilibrium states, generation times, reproductive values, and related quantities, often supported by empirical analyses using plant and animal data. A continuous-time analogue, the Path Integral Model (Oizumi and Takada, 2013; Oizumi, 2022b), derives population-level dynamics from stochastic models of individual life histories and can incorporate optimization principles such as $r/K$-selection (Oizumi et al., 2016). It can be viewed as a variant of the multi-state McKendrick equation, which generalizes the McKendrick–von Foerster PDE by including state variables beyond age.
@@
Mathematically, these quantities are governed by the spectral radius of the associated positive integral operator and its eigenfunctions, which are often characterized by a Fredholm integral equation of the second kind. Fredholm’s original solution (Fredholm, 1903), expressed via the Fredholm determinant, is foundational in functional analysis, but determinant-based formulations can be computationally expensive and may provide limited biological transparency in applied settings. For practical inference, many studies adopt quadrature-based matrix approximations (Ellner and Rees, 2006; White et al., 2016) that discretize the state space; however, discretization requires care because it may not preserve qualitative properties of the original continuous model.
In demographic modeling, a classical alternative to determinant calculations is the Euler–Lotka equation (Euler, 1760; Sharpe and Lotka, 1911), which is equivalent to the characteristic equation for the intrinsic growth rate but is more tractable and biologically transparent, relating directly to net reproduction and generation-time distributions (Inaba, 2017). Recently, Oizumi et al. (2022a) extended a classical Markov chain idea (Chung, 1960) to irreducible nonnegative matrices, expressing eigenvectors associated with the Frobenius root explicitly in terms of matrix entries and taboo probabilities.
@@
Our key device is a reference-point operator (a rank-one operator at the kernel level) constructed from a fixed reference pair $(x_{0},y_{0})$. This operator algebraically implements a Markov-chain–inspired “taboo” decomposition at the reference point in a way that remains well posed for continuous-state kernels, and it leads to a family of reference-point iterates $\Gamma_{n}$ generated by $(\mathbf{K}-\mathbf{P})$. The resulting Neumann-series/resolvent expansion replaces determinant expressions by sums over iterated kernels, so each term admits a direct genealogical reading as a multi-generation contribution. Moreover, the combinatorics of the iterates admit a closed representation in terms of ordinary partial Bell polynomials, yielding an explicit expression of $\Gamma_{n}$ and the associated reference-point moments. These moments in turn provide a characteristic equation in an Euler–Lotka-like form, which is well suited to positivity-based analysis and biological interpretation.
@@ 2.1. Kernel space, kernels, and iterates
$\|F\|_{\infty,U}:=\sup_{(x,y)\in U}|F(x,y)|,\qquad\mathcal{X}_{\star}:=\{F\in\mathcal{X}:\ \|F\|_{\infty,U}<\infty\},\qquad\|F\|_{\mathcal{X}_{\star}}:=\|F\|_{\mathcal{X}}+\|F\|_{\infty,U}.$
@@
(4) $K^{(1)}(x,y):=K(x,y),\qquad K^{(n+1)}(x,y):=(\mathbf{K}K^{(n)})(x,y)=\int_{\Omega^{d}}K(x,\xi)\,K^{(n)}(\xi,y)\,d\xi,\quad n\geq 1.$
@@
$\mathbb{K}_{2}:=\left\{K\in\mathbb{K}\,\middle|\,\int_{\Omega^{d}\times\Omega^{d}}|K(x,y)|^{2}\,dx\,dy<\infty\,and\int_{\Omega^{d}}|K(x,x)|^{2}\,dx<\infty\right\},$
@@ 2.2. Non-Hilbert–Schmidt solution and its property
(7) $\Gamma_{n}(x,y,x_{0},y_{0})=K^{(n)}(x,y)+\sum_{\ell=1}^{n-1}(-1)^{\ell}\sum_{k=\ell}^{n-1}K^{(n-k)}(x,y)\,\widehat{B}_{k,\ell}\!\Big(K^{(1)},K^{(2)},\dots,K^{(k)};x_{0},y_{0}\Big).$
@@
Following the original method of Fredholm (Fredholm, 1903), which is based on a formal series expansion of the kernel through successive iterations, we expand the solution $w$ into a series normalized by the nonzero constant $c_{0}(x_{0},y_{0})$:
@@
$\Gamma_{n+1}(x,y;x_{0},y_{0})$ $=\int_{\Omega^{d}}K(x,\xi)\Gamma_{n}(\xi,y;x_{0},y_{0})\,d\xi-K(x,y)\Gamma_{n}(x_{0},y_{0};x_{0},y_{0}),\qquad n\geq 1.$
Step 3.
@@
$w(x,y,x_{0},y_{0};\lambda)=\frac{1}{\lambda}c_{0}(x_{0},y_{0})\left[\left(\mathbf{I}-\frac{1}{\lambda}\mathbf{A})\right)^{-1}K\right](x,y).$
@@
for the one-dimensional linear subspace generated by $v$ (and similarly $\mathrm{span}\{v_{1},\dots,v_{m}\}:=\{\sum_{i=1}^{m}\alpha_{i}v_{i}:\alpha_{i}\in\mathbb{R}\}$).
@@
(36) $\Gamma_{n+1}^{*}(y_{0},x_{0};y,x)=\int_{\Omega^{d}}\Gamma_{n}^{*}(y_{0},x_{0};y,\xi)\,K(\xi,x)\,d\xi-\Gamma_{n}^{*}(y_{0},x_{0};y_{0},x_{0})\,K^{\top}(x,y),\qquad n\geq 1.$
@@
(37) $v(y,x;\lambda)=\frac{c_{1}(x_{0},y_{0})}{\lambda}\left[\left(\mathbf{I}-\frac{1}{\lambda}(\mathbf{K}^{*}-\mathbf{P}^{*})\right)^{-1}K^{\top}\right](x,y)=c_{1}(x_{0},y_{0})\sum_{n=1}^{\infty}\frac{1}{\lambda^{n}}\,\Gamma_{n}^{*}(y_{0},x_{0};y,x),$
@@
Step 1 (Generating-function identity). Let $\rho(\mathbf{K})$ be the spectral radius of $\mathbf{K}.$ Then for every $\lambda>\rho(\mathbf{K})$, the resolvent exists and we have
@@ 3. Application to simple integral projection models at discrete time
In this section, we investigate the eigenvalue problem for the simplest IPM, using the eigenfunction representation provided by Theorem 2.6. This representation also offers a Markovian and biological interpretation of the system.
Let $P_{t}(x)$ denote the cohort at state $x$ at time $t$. The simplest IPM considered here is given by
(39) $P_{t+1}\left(x\right)=\int_{\Omega^{d}}K\left(x,y\right)P_{t}\left(y\right)\,dy,\quad P_{0}\left(x\right)\in L^{1}\left(\Omega^{d}\right),\quad P_{0}\left(x\right)\geq 0,$
where the integral kernel $K\left(\cdot\right)$ is defined in Theorem 2.6.
Before constructing the eigenvalue problem for this IPM, we introduce one convention. The representation used in Theorem 2.6 involves three points not directly related to the integral operator, namely $y$, $x_{0}$, and $y_{0}$. These points can be chosen freely provided that the chosen reference pair satisfies the condition of Lemma 2.16. However, in consideration of the usefulness of biological interpretation, which will be discussed later, we therefore impose the convention $y=x_{0}=y_{0}$ and choose the reference point $y_{0}$ so that the associated pair $(y_{0},y_{0})$ satisfies Lemma 2.16. Based on these assumptions, we will omit variables from this chapter as follows:
@@ 3.2. Reinterpretation of eigensystems by analogy with Markov chains
In understanding the sequence appearing in (40), the theory of Markov chains offers particularly valuable suggestions. In a Markov chain whose transition probability from state $j$ to state $i$ is denoted by $p_{ij}\geq 0$, the following quantity, called the taboo probability, is known (Chung, 1960).
@@
This relation does not require the matrix to be stochastic; an irreducible nonnegative matrix similarly yields an eigenvector corresponding to its Frobenius root (Oizumi et al., 2022a). Focusing on the right-hand side of (51), we see that the first term sums over all paths from every state $k$ to state $i$ at the previous step, while the second term subtracts the contribution of the paths that pass through state $j$. Then, by replacing the sum with an integral in (51), we observe that the resulting relation resembles the difference equation (42) that the coefficients of each power of $\lambda_{0}$ in (40) should satisfy.
@@ 3.3. Initial population dependence and expected contribution steps
$=\langle\Gamma_{m-1}^{*},\Gamma_{n+1}\rangle_{x}+\int_{\Omega^{d}}\Gamma_{m-1}^{*}(y_{0},\eta)\,K(\eta,y_{0})\,\Gamma_{n}\,d\eta-\Gamma_{m-1}\int_{\Omega^{d}}K(y_{0},x)\,\Gamma_{n}(x,y_{0})\,dx$
@@
(57) $\langle v_{0},w_{0}\rangle_{x}$ $=\int_{\Omega^{d}}v_{0}(y_{0},x)\,w_{0}(x,y_{0})\,dx$
(58) $=c_{v}(y_{0})\,c_{w}(y_{0})\int_{\Omega^{d}}\left(\sum_{m=1}^{\infty}\frac{\Gamma_{m}^{*}(y_{0},x)}{\lambda_{0}^{\,m}}\right)\left(\sum_{n=1}^{\infty}\frac{\Gamma_{n}(x,y_{0})}{\lambda_{0}^{\,n}}\right)\,dx$
(59) $=c_{v}(y_{0})\,c_{w}(y_{0})\sum_{m=1}^{\infty}\sum_{n=1}^{\infty}\frac{1}{\lambda_{0}^{\,m+n}}\int_{\Omega^{d}}\Gamma_{m}^{*}(y_{0},x)\,\Gamma_{n}(x,y_{0})\,dx$
[…]
@@
The quantity $T_{y_{0}}$ is referred to as the type reproduction number (TRN) (Heesterbeek and Roberts, 2007; Inaba, 2013) at state $y_{0}$. By analogy with Markov chains, it aggregates, over all $n$, the total contribution from individuals originating in state $y$ who either return to $y_{0}$ or produce descendants that reach $y_{0}$ for the first time at step $n$. From a measure-theoretic viewpoint, a single point in $\mathbb{R}^{d}$ has zero recurrence measure; thus this interpretation is merely heuristic. Nevertheless, given the meaning of the quantity and its relation to the dominant spectral value $\lambda_{0}$, it is natural to regard $T_{y_{0}}$ as the type reproduction number associated with a single state.
@@ 4.1. Assumptions of Multi-state McKendrick equation
The Laplace transform is a standard tool for analyzing the asymptotics of renewal equations with integral operators (see (Inaba, 2017)). For $g\in L^{1}(0,\infty)$ and $r\in\mathbb{C}$, set
@@
The series in (82) generally diverges (Feller, 1941); in many cases, solutions cannot even be expressed by a series (Verduyn Lunel, 1989, 1990). However, if $\mathbf{\Psi}(a)$ is generated by a Hilbert–Schmidt kernel, the existence and uniqueness of solutions to (75) and the simplicity of the leading eigenvalue $r_{0}$ follow from Mode (1971).

### v3

@@ 1. Introduction
The Integral Projection Model (IPM) is a mathematical framework for describing transitions in populations with continuous state variables Ellner and Rees (2006). In discrete-time settings, IPMs have been developed and applied widely in ecology and demography. For example, Coulson et al. (2010) studied trait evolution, while Coulson and Tuljapurkar (2008) analyzed asymptotic population dynamics under environmental variability. These and other studies White et al. (2016); Nicolè et al. (2011); Doak et al. (2021); Merow et al. (2014); Elderd and Miller (2016) have linked spectral theory to equilibrium states, generation times, reproductive values, and related quantities, often supported by empirical analyses using plant and animal data. A continuous-time analogue, the Path Integral Model Oizumi and Takada (2013); Oizumi (2022a), derives population-level dynamics from stochastic models of individual life histories and can incorporate optimization principles such as $r/K$-selection Oizumi et al. (2016). It can be viewed as a variant of the multi-state McKendrick equation, which generalizes the McKendrick–von Foerster PDE McKendrick (1925); von Foerster (1959) by including state variables beyond age.
@@
Mathematically, these quantities are governed by the spectral radius of the associated positive integral operator and its eigenfunctions, which are often characterized by a Fredholm integral equation of the second kind. Fredholm’s original solution Fredholm (1903), expressed via the Fredholm determinant, is foundational in functional analysis, but determinant-based formulations can be computationally expensive and may provide limited biological transparency in applied settings. For practical inference, many studies adopt quadrature-based matrix approximations Ellner and Rees (2006); White et al. (2016) that discretize the state space; however, discretization requires care because it may not preserve qualitative properties of the original continuous model.
In demographic modeling, a classical alternative to determinant calculations is the Euler–Lotka equation Euler (1760); Sharpe and Lotka (1911), which is equivalent to the characteristic equation for the intrinsic growth rate but is more tractable and biologically transparent, relating directly to net reproduction and generation-time distributions Inaba (2017). Recently, Oizumi et al. (2022b) extended a classical Markov chain idea Chung (1960) to irreducible nonnegative matrices, expressing eigenvectors associated with the Frobenius root explicitly in terms of matrix entries and taboo probabilities.
@@
Our key device is a reference-point operator (a rank-one operator at the kernel level) constructed from a fixed reference pair $(x_{0},y_{0})$. This operator algebraically implements a Markov-chain–inspired “taboo” decomposition at the reference point in a way that remains well posed for continuous-state kernels, and it leads to a family of reference-point iterates $\Gamma_{n}$ generated by $(\mathbf{K}-\mathbf{P})$. The resulting Neumann-series/resolvent expansion Neumann (1877) replaces determinant expressions by sums over iterated kernels, so each term admits a direct genealogical reading as a multi-generation contribution. Moreover, the combinatorics of the iterates admit a closed representation in terms of ordinary partial Bell polynomials, yielding an explicit expression of $\Gamma_{n}$ and the associated reference-point moments. These moments in turn provide a characteristic equation in an Euler–Lotka-like form, which is well suited to positivity-based analysis and biological interpretation.
@@ 2.1. Kernel space, kernels, and iterates
$\|F\|_{\infty,U}$ $:=\sup_{(x,y)\in U}|F(x,y)|,\quad\mathcal{X}_{\star}:=\{F\in\mathcal{X}:\ \|F\|_{\infty,U}<\infty\},$
$\|F\|_{\mathcal{X}_{\star}}$ $:=\|F\|_{\mathcal{X}}+\|F\|_{\infty,U}.$
@@
(4) $K^{(1)}(x,y):=K(x,y),\qquad K^{(n+1)}(x,y)$ $:=(\mathbf{K}K^{(n)})(x,y)$
$=\int_{\Omega^{d}}K(x,\xi)\,K^{(n)}(\xi,y)\,d\xi,\quad n\geq 1.$
@@
$\mathbb{K}_{2}$={} $\{K\in\mathbb{K}\mid\int_{\Omega^{d}\times\Omega^{d}}|K(x,y)|^{2}\,dx\,dy<\infty$
$\text{and }\int_{\Omega^{d}}|K(x,x)|^{2}\,dx<\infty\},$
@@ 2.2. Non-Hilbert–Schmidt solution and its property
(7) $\Gamma_{n}(x,y,x_{0},y_{0})=$ $K^{(n)}(x,y)$
$+\sum_{\ell=1}^{n-1}(-1)^{\ell}\sum_{k=\ell}^{n-1}K^{(n-k)}(x,y)\,\widehat{B}_{k,\ell}\!\Big(K^{(1)},K^{(2)},\dots,K^{(k)};x_{0},y_{0}\Big).$
@@
Following the original method of Fredholm Fredholm (1903), which is based on a formal series expansion of the kernel through successive iterations, we expand the solution $w$ into a series normalized by the nonzero constant $c_{0}(x_{0},y_{0})$:
@@
$\Gamma_{n+1}(x,y;x_{0},y_{0})$ $=\int_{\Omega^{d}}K(x,\xi)\Gamma_{n}(\xi,y;x_{0},y_{0})\,d\xi-K(x,y)\Gamma_{n}(x_{0},y_{0};x_{0},y_{0}),\quad n\geq 1.$
Step 3:
@@
$w(x,y,x_{0},y_{0};\lambda)=\frac{1}{\lambda}c_{0}(x_{0},y_{0})\left[\left(\mathbf{I}-\frac{1}{\lambda}\mathbf{A}\right)^{-1}K\right](x,y).$
@@
for the one-dimensional linear subspace generated by $v$ (and similarly $\mathrm{span}$
$\{v_{1},\dots,v_{m}\}:=\{\sum_{i=1}^{m}\alpha_{i}v_{i}:\alpha_{i}\in\mathbb{R}\}$).
@@
(36) $\Gamma_{n+1}^{*}(y_{0},x_{0};y,x)=\int_{\Omega^{d}}\Gamma_{n}^{*}(y_{0},x_{0};y,\xi)\,K(\xi,x)\,d\xi-\Gamma_{n}^{*}(y_{0},x_{0};y_{0},x_{0})\,K^{\top}(x,y),$
$n\geq 1.$
@@
$v(y,x;\lambda)$ $=\frac{c_{1}(x_{0},y_{0})}{\lambda}\left[\left(\mathbf{I}-\frac{1}{\lambda}(\mathbf{K}^{*}-\mathbf{P}^{*})\right)^{-1}K^{\top}\right](x,y)$
$=c_{1}(x_{0},y_{0})\sum_{n=1}^{\infty}\frac{1}{\lambda^{n}}\,\Gamma_{n}^{*}(y_{0},x_{0};y,x),$
@@
Step 1: Generating-function identity. Let $\rho(\mathbf{K})$ be the spectral radius of $\mathbf{K}.$ Then for every $\lambda>\rho(\mathbf{K})$, the resolvent exists and we have
@@ 3. Application to simple integral projection models at discrete time
In this section, we investigate the eigenvalue problem for the simplest discrete-time IPM, using the eigenfunction representation provided by Theorem 2.6. This representation also yields a Markovian viewpoint and facilitates biological interpretation.
Let $P_{t}(x)$ denote the cohort density at state $x$ at time $t$. We consider the IPM
(38) $P_{t+1}(x)=\int_{\Omega^{d}}K(x,y)\,P_{t}(y)\,dy,\qquad P_{0}\in L^{1}(\Omega^{d}),\quad P_{0}\geq 0,$
where the kernel $K\in\mathbb{K}$ is assumed to admit a diagonal reference point.
Assumption 3.1.
There exists $y_{0}\in\Omega^{d}$ such that
$K(y_{0},y_{0})>N_{0}-\rho(\mathbf{K}).$
This assumption guarantees that the reference-point series,
$\sum_{n\geq 1}\lambda^{-n}\Gamma_{n}(\cdot,\cdot;y_{0},y_{0}),$
converges at $\lambda=\rho(\mathbf{K})$, even when evaluated at the diagonal point $(y_{0},y_{0})$.
Assumption 3.1 is imposed solely to keep the subsequent results biologically interpretable; it is not mathematically essential for the analysis that follows. Without it, the same arguments go through at the expense of slightly more cumbersome notation and formulas. Nevertheless, this assumption is convenient for presenting a clean mathematical structure—in particular, it allows us to state and prove Theorem 3.4 later in the paper in a transparent form.
Remark 3.2 (A biological interpretation of Assumption 3.1).
This condition is natural in many biological settings, for instance for long-lived organisms with slow growth such as trees, for species whose adult size is essentially fixed (as is often the case in mammals), and for organisms exhibiting strong site fidelity—that is, individuals that remain in the same habitat or patch with little movement.
Before formulating the eigenvalue problem for (38), we introduce a convention regarding the reference points appearing in Theorem 2.6. The representation in Theorem 2.6 involves three auxiliary points, denoted by $y$, $x_{0}$, and $y_{0}$, which may be chosen freely as long as the associated reference pair satisfies the condition of Lemma 2.16. However, in view of the biological interpretation discussed below, we impose the convention
$y=x_{0}=y_{0}$
and choose the reference point $y_{0}$ so that the associated pair $(y_{0},y_{0})$ satisfies Lemma 2.16. Since $\Omega^{d}\subseteq\mathbb{R}^{d}$ and $K$ is continuous, the diagonal map $y\mapsto K(y,y)$ is continuous; hence, in view of Assumption 3.1, we assume that $y_{0}$ can be chosen so that $K(y_{0},y_{0})>N_{0}-\rho(\mathbf{K})$.
Based on these assumptions, we will omit variables from this chapter as follows:
@@ 3.2. Reinterpretation of eigensystems by analogy with Markov chains
In understanding the sequence appearing in (39), the theory of Markov chains offers particularly valuable suggestions. In a Markov chain whose transition probability from state $j$ to state $i$ is denoted by $p_{ij}\geq 0$, the following quantity, called the taboo probability, is known Chung (1960).
@@
This relation does not require the matrix to be stochastic; an irreducible nonnegative matrix similarly yields an eigenvector corresponding to its Frobenius root Oizumi et al. (2022b). Focusing on the right-hand side of (50), we see that the first term sums over all paths from every state $k$ to state $i$ at the previous step, while the second term subtracts the contribution of the paths that pass through state $j$. Then, by replacing the sum with an integral in (50), we observe that the resulting relation resembles the difference equation (41) that the coefficients of each power of $\lambda_{0}$ in (39) should satisfy.
@@ 3.3. Initial population dependence and expected contribution steps
$=\langle\Gamma_{m-1}^{*},\Gamma_{n+1}\rangle_{x}+\int_{\Omega^{d}}\Gamma_{m-1}^{*}(y_{0},\eta)\,K(\eta,y_{0})\,\Gamma_{n}\,d\eta$
$\quad-\Gamma_{m-1}\int_{\Omega^{d}}K(y_{0},x)\,\Gamma_{n}(x,y_{0})\,dx$
@@
$\langle v_{0},w_{0}\rangle_{x}$ $=\int_{\Omega^{d}}v_{0}(y_{0},x)\,w_{0}(x,y_{0})\,dx$
$=c_{v}(y_{0})\,c_{w}(y_{0})\int_{\Omega^{d}}\left(\sum_{m=1}^{\infty}\frac{\Gamma_{m}^{*}(y_{0},x)}{\lambda_{0}^{\,m}}\right)\left(\sum_{n=1}^{\infty}\frac{\Gamma_{n}(x,y_{0})}{\lambda_{0}^{\,n}}\right)\,dx$
$=c_{v}(y_{0})\,c_{w}(y_{0})\sum_{m=1}^{\infty}\sum_{n=1}^{\infty}\frac{1}{\lambda_{0}^{\,m+n}}\int_{\Omega^{d}}\Gamma_{m}^{*}(y_{0},x)\,\Gamma_{n}(x,y_{0})\,dx$
$=c_{v}(y_{0})\,c_{w}(y_{0})\sum_{m=1}^{\infty}\sum_{n=1}^{\infty}\frac{1}{\lambda_{0}^{\,m+n}}\left(\Gamma_{m+n}(y_{0},y_{0})+\Gamma_{m}(y_{0},y_{0})\,\Gamma_{n}(y_{0},y_{0})\right)$
(by Theorem 3.4)
@@
The quantity $T_{y_{0}}$ is referred to as the type reproduction number (TRN) Heesterbeek and Roberts (2007); Inaba (2013) at state $y_{0}$. By analogy with Markov chains, it aggregates, over all $n$, the total contribution from individuals originating in state $y$ who either return to $y_{0}$ or produce descendants that reach $y_{0}$ for the first time at step $n$. From a measure-theoretic viewpoint, a single point in $\mathbb{R}^{d}$ has zero recurrence measure; thus this interpretation is merely heuristic. Nevertheless, given the meaning of the quantity and its relation to the dominant spectral value $\lambda_{0}$, it is natural to regard $T_{y_{0}}$ as the type reproduction number associated with a single state.
@@ 4.1. Assumptions of Multi-state McKendrick equation
The Laplace transform is a standard tool for analyzing the asymptotics of renewal equations with integral operators (see Inaba (2017)). For $g\in L^{1}(0,\infty)$ and $r\in\mathbb{C}$, set
@@
The series in (78) generally diverges Feller (1941); in many cases, solutions cannot even be expressed by a series Verduyn Lunel (1989, 1990). However, if $\mathbf{\Psi}(a)$ is generated by a Hilbert–Schmidt kernel, the existence and uniqueness of solutions to (71) and the simplicity of the leading eigenvalue $r_{0}$ follow from Mode (1971).